### v1

# Non-reciprocal neutral ferroelectric domain walls in BiFeO3

Abstract
This paper analyzes a peculiar phenomenon of non-reciprocal domain wall pairs and illustrate the implications in ab-initio-based atomistic computational experiments with ($\bar{1}\bar{1}2$)-oriented planar R180 domain walls within the canonical multiferroic ferroelectric crystal of BiFeO3. Results show that parallel walls on the opposite sides of a given domain within a simply twinned lamellar domain structure can have considerably different polarization and oxygen octahedra tilt profiles, different thickness and energy densities. The spontaneous formation of zigzag walls and triangular domains suggests that these domain walls are actually the lowest energy planar R180 walls in pure insulating BiFeO3 crystals and films.
Ferroelectric BiFeO3, Zigzag domain walls
pacs:
42.50.Vk, 78.30.-j, 63
Domain structure is a natural manifestation of symmetry breaking phase transitions in crystals. One of the simplest but still very frequent domain structure is the simple lamellar twinning, consisting in alternation of only two distinct domains states, $A$ and $B$, separated by parallel planar domain walls. Often, the two subsequent walls $A|B$ and $B|A$ are equivalent by symmetry reasons. However, it is also possible that the symmetry relationship between A||||B and B||||A is absent. Then, $A|B$ and $B|A$ are physically non-equivalent walls. Such non-reciprocal walls were earlier also referred as irreversible onesJanovec .
The non-reciprocity applies to all nominally charged ferroelectric domain walls. Such walls show a mismatch in the normal component of the spontaneous polarization in the adjacent domains. In lamellar case, the two subsequent charged walls can be denoted head-to-head (HH) and tail-to-tail (TT) wall, respectively. The HH and TT walls differ by the sign of the interfacial charge and thus the absence of symmetry relation is obvious. For example, it was reported that the electric conductivity of HH and TT 90 walls in barium titanate differs by several orders of magnitude.
In this paper, we are addressing the question whether and when one can encounter the manifestations of a much less obvious non-reciprocity of nominally neutral ferroelectric walls. In particular, we explore the case of the experimentally reported but so far little investigated $(\bar{1}2\bar{1})$ oriented 180 walls in BiFeO3 and argue that these walls are indeed showing remarkable non-reciprocal effects.
BiFeO3 is a well-studied multiferroic material with coexisting ferroelectric, ferroelastic, antiferrodistortive and antiferromagnetic orders and promising applications in spintronics, nonvolatile memory devices, catalysis and sensors wang_2020 ; chaudron_2024 ; YangAbove ; Amdouni . At room temperature, it has rhombohedral $R3c$ symmetry with a spontaneous polarization between $0.6-1.0~{}C/m^{2}$ and with oxygen octahedra alternatively tilted by about $13.8^{\circ}~{}$ around the polarization axis  lebeugle_2007 ; Catalan09 . The conventional hexagonal unit cell is depicted in Fig. 1. The parent cubic perovskite $m\bar{3}m$ reference phase gives rise to eight equivalent polarization states. Corresponding ferroelectric domain walls are primarily distinguished as ${}^{\circ}~{}$, ${}^{\circ}~{}$ and ${}^{\circ}~{}$ ones according to the angles between the polarization in the adjacent domains. Details structure and properties of domain walls in BiFeO3 have been subject of many experimentalFarokhipoor ; Sando14 ; Bhatnagar and theoretical studiesDieguez2013 ; Ren2013 ; ourPRL . The benchmark systematic first-principles study of Ref. Dieguez2013, explored almost all types of mechanically and electrically compatible (neutral) domain walls. However, non-reciprocal walls were not addressed there. The present work fills this knowledge gap and employ an atomistic model to demonstrate and explain several remarkable non-reciprocal domain wall effects with specific implications for this important model ferroelectric material.
To obtain 180 domain walls we used an interatomic potential with parameters fitted from ab-initio calculations Graf2014 ; Graf2014b that has been also employed in Refs. Graf2015, ; Graf2016, ; ourPRL, . The potential utilizes the framework of the shell model, representing each atom in the system with a core and a shell, mimicking the atomic polarizability and with the anharmonic core-shell interaction is particularly well-suited in reproducing ferroelectric behavior  Tinte2004 . The low-temperature multidomain configuration was stabilized using a supercell made of 20$\times$20$\times$2 rhombohedral BiFeO3 cells in the hexagonal setting (i.e. 24000 atoms) subjected to periodic boundary conditions.
The structural relaxation was achieved using classical molecular dynamics simulations under constant stress and temperature conditions. A temperature of 1 K was chosen to facilitate effective relaxation while introducing a minimal amount of thermal noise. The simulation began with a 10 ps thermalization, followed by an additional 10 ps during which the average configuration was computed. The molecular dynamics calculations were performed using the DL_POLY software  Todorov2006 . Local dipole moments were evaluated for each 5-atom perovskite cell centered in the Bi atoms, based on the positions and charges of cores and shells. For a given cell, one Bi atom with its surrounding twelve O atoms (with a weight of 0.25) and eight coordinating Fe atoms (0.125 weight) were considered.
Figure 1: BiFeO3 structure and the orientation of the investigated R180($\bar{1}2\bar{1}$) domain walls. (a) hexagonal cell and the corresponding lattice vectors. Purple spheres stand for Bi, red spheres for O, dark yellow for Fe. (b) Schema of the simulation box with two domain walls, parallel to the plane defined by $\bf a_{\rm h}$ and $\bf b_{\rm h}$ lattice vectors. Central (orange) domain stripe has polarization in a positive sense of $z^{\prime}$ axis, cyan domains have polarization in the opposite sense. Adopted cartesian coordinate system $x^{\prime}$, $y^{\prime}$ and $z^{\prime}$ is also indicated.
The present study focuses on ${}^{\circ}~{}$ domain walls. Keeping the usual notation where the Miller indices $h,k,l$ refer to the parent cubic structure. We choose two antiparallel domain states with polarization along $[111]$ and $[\bar{1}\bar{1}\bar{1}]$. The mechanically and electrically compatible ${}^{\circ}~{}$ domain walls connecting these two states coincide with planes parallel to the $[111]$ direction, that we also identify with positive sense of $z^{\prime}$ axis (see Fig. 1a). There are two high-symmetry options for such a plane; $(1\bar{1}0)$-type planes, parallel to the mirror symmetry plane of the parent $m\bar{3}m$ reference phase, and ($\bar{1}\bar{1}2$)-type planes, passing through the intersecting 3-fold and 2-fold rotational axes of the parent $m\bar{3}m$ reference phase. Most of the studies have considered only the former type, but as we shall argue, the ($\bar{1}\bar{1}2$)-type walls are even more important for BiFeO3.
The desired ($\bar{1}\bar{1}2$) domain wall orientation was selected by imposing the suitable step-like distribution of the initial Bi ion off-centering displacements in our simulation box. In order profit from the computational advantages of the periodic boundary conditions, one has to introduce two such domain walls, as indicated in Fig. 1b.
After the relaxation of the structure, most of the volume is in the equilibrium single domain structure with the spontaneous polarization either parallel or antiparallel to the sense of the $z^{\prime}$ axis. For the detailed inspection of the domain walls, polarization of each unit cell was projected on the rotated coordinate system, where $s\parallel[\bar{1}\bar{1}2]$ is perpendicular to the wall, $z^{\prime}\parallel[111]$ is along the polarization, and $t\parallel{\bf b_{\rm h}}$ is perpendicular to the both $s$ and $z^{\prime}$.
Resulting polarization profiles in the vicinity of domain wall DW1 and DW2 are shown in Fig. 2a,b. Analogous profiles of the staggered oxygen octahedra tilt rotations $R_{i}$ are shown in Fig. 2cd. As expected, in both DW1 and DW2 the $P_{\rm s}$ and $P_{\rm t}$ components interpolate between zero values, while the $P_{\rm z^{\prime}}$ component is switching between the nominal spontaneous polarization values. The absence of switching of the $R_{\rm i}$ components confirms that neither DW1 nor DW2 coincide with an anti-phase boundary of the underlying staggered tilt pattern and thus both walls can be considered as purely ferroelectric domain walls.
While the profiles of $P_{\rm s}$, $P_{\rm z^{\prime}}$, $R_{\rm s}$, and $R_{\rm z^{\prime}}$ components are same at any line passing through a given wall, the values of $P_{\rm t}$ and $R_{\rm t}$ alternate between two mutually opposite values when comparing lines related by the fractional translation vector ${\bf c_{\rm h}}/2$. It means that the Bloch polarization component averaged over the two sites related by ${\bf c_{\rm h}}/2$ is zero, and at the same time, there is a staggered (antiferroelectric) Bloch component within the wall. In case of the tilt profile, the double signs of $R_{\rm t}$ implies that the Bloch component of the staggered oxygen octahedra rotation averaged over the two octahedra related by ${\bf c_{\rm h}}/2$ is zero, while both octahedra are slightly rotated in the same sense around the $t$-axis.
However, there are also important differences between DW1 and DW2. First, the DW1 wall has thickness of about 0.7 nm while DW2 thickness is about 0.35 nm Second, the magnitude of the tilt modulations are much stronger in DW1 and moreover, the signs of the $R_{\rm s}$ are opposite in DW1 and DW2. A mode detailed analysis of the stuctural profiles within a continuous Ginzburg-Landau theory is a subject of a separate report.
Figure 2: Profiles of the components polarization $P_{i}$ (a,b) and oxygen octahedra tilts $R_{i}$ (c,d) across the DW1 (a,c) and DW2 (b,d) domain walls defined in Figure 1b. The rotated cartesian coordinate system $s,t,z^{\prime}$ is used.
In order to verify that the resulting profiles shown in Fig. 2 are the lowest energy domain wall configurations, we have slightly varied the initial positions of the walls, size of the simulation box and we also annealed at higher temperatures. In overall, the domain characteristics described proved to be robust and representative ones.
Nevertheless, a very interesting phenomenon was encountered in the course of gradual temperature increase in steps of 20 K up to 300 K. This computer experiment, documented in Fig. 3, has been performed in the supercell made of $20\times 20\times 8$ rhombohedral hexagonal cells (i.e. 96000 atoms). The molecular dynamics trajectory for each temperature began with a 10 ps thermalization, followed by an additional 10 ps for computing temporal averages from snapshots recorded every 0.05 ps. The final configuration of the previous temperature was used as the starting point for the next temperature.
The time-averaged polarization map at $z^{\prime}=0$ was practically same till $T=120$ K (see Fig. 3a). However, when the temperature reached $T=140$ K, the DW2 started to display significant fluctuations. The original planar wall changed to an irregularly corrugated shape and the $P_{\rm t}$ components started to be disordered. The slow fluctuations are apparent even in the 10 ps averaged polarization maps (see Fig. 3b). Finally, at $T=240$ K, the habitus of DW2 abruptly transformed into a zigzag wall with ${}^{\circ}~{}$ folds. This form of DW2 then persists till $T=300$ K (see Fig. 3c) and it remains there even when they system is cooled back down to $T=1$ K.
Figure 3: Calculated polarization maps showing evolution of the 180 domain walls while increasing the temperature to (a) $100$ K, (b)$140$ K and (c) $300$ K. Arrows show the non-negligible in-plane polarization components, the color shows the out-of-plane polarization component.
Finally, we also tried to stabilize small diameter columnar nanodomain of inverted polarization, still with neutral domain walls, but otherwise without any constraint from the periodicity of boundary conditions in the $x^{\prime}y^{\prime}$ plane. Obviously, small wall segments or the curved domain walls are more strongly interacting but they are also more free to change their position or orientation. For a $P_{\rm z^{\prime}}>0$ nanodomain, the only encountered stable configurations had a cross section of an equilateral triangle, oriented as shown in Fig. 4. We could see that one of the sides of this right triangular prism is always parallel the walls of Fig. 1b and that the wall profiles correspond to that of DW1 in Fig. 2ac. Moreover, we never seen $(1\bar{1}0$-type walls suggesting that DW1 is actually the lowest energy 180 domain wall.
Summarizing, we have shown that the DW2 differs from DW1 the thickness, amplitudes of the secondary order parameters in the walls and also by the respective stability with respect to temperature and domain wall interactions. The ensemble of all these findings demonstrates that DW1 and DW2 are not related by any symmetry operation.
Figure 4: Simulations of triangular nanodomains. (a)Polarization map in of the relaxed structure of a triangular nanodomain. (b) The excess energy due to the domain wall as a function of the lateral extension of the nanodomain and the linear fit to the data.
For a more precise discussion, let us introduce a more instructive domain wall symbols $A|{\bf{n}}\rangle B]$. Here $A$ and $B$ stands for domain states while the oriented domain wall normal $\bf{n}$ points in the real space towards the domain that is ascribed to the state of the right-hand side of the symbol, as it is mnemonically indicated by the embedded ”ket” symbol.
Therefore, using the already introduced pseudocubic Miller indices for spontaneous polarization directions and oriented domain wall normals, our DW1 of Fig. 1b can be described as either $[111|1\bar{2}1\rangle\bar{1}\bar{1}\bar{1}]$ or $[\bar{1}\bar{1}\bar{1}|\bar{1}2\bar{1}\rangle 111]$, while DW2 can be only described as $[111|\bar{1}2\bar{1}\rangle\bar{1}\bar{1}\bar{1}]$ or $[\bar{1}\bar{1}\bar{1}|1\bar{2}1\rangle 111]$.
Symmetry of planar domain walls was systematically developed by JanovecJanovec . The basic assumption is that specific symmetries are described by certain subgroups of the parent phase crystallographic group $G$. The proper symmetry of the $[A|{hkl}\rangle B]$ wall is described by a group $T_{AB}|{hkl}\rangle$, composed of those parent phase symmetry group operations, that results in an indistinguishable domain wall geometry. It is a crystallographic group with two-dimensional translational symmetry. The highest possible symmetry $T_{AB}(hkl)$ can be determined considering the orientation of the domain wall plane and its centering in the crystal structure. In the present case, the $2_{\rm t}/m_{\rm t}$ factor group of $T_{AB}(hkl)$ contains the identity, inversion $i$, mirror $m_{\rm t}$ perpendicular to $t$ axis and 2-fold rotation $2_{\rm t}$ parallel to the $t$ axis. Both planar DW1 and planar DW2 walls agree with this symmetry within the precision of our numerical calculations.
The non-reciprocity is related to the symmetry group of domain wall pair $\bar{J}_{AB}|(hkl)$, which collects all parent phase symmetry operations that preserve or exchange the domain states and preserve the domain wall plane. In principle, this is supergroup of $T_{AB}(hkl)$ because it can also contain operations that interchange the sides of the wall but do not alter the domain state (pure side-reversal operations) or operations that reverse the domain states but do not interchange sides of the wall (pure state-reversal operations). Existence of pure side-reversal or pure state-reversal symmetry implies the reciprocity (reversibility) of the wall. In the case of our wall, no such symmetry operations exists and $\bar{J}_{AB}|(hkl)=T_{AB}(hkl)$.
Finally, the group of the unordered domain pair, $J_{AB}$, involves all operations maintaining or interchanging the two domain states. In our case the $\bar{3}m$ factor group of $J_{AB}$ is a nontrivial supergroup of the factor group of $\bar{J}_{AB}|(hkl)$ what implies that there are the walls equivalent to DW1 that joins the same two domain states. It can be seen that these are the $[111|\bar{2}11\rangle\bar{1}\bar{1}\bar{1}]=[\bar{1}\bar{1}\bar{1}|\bar{1}2\bar{1}\rangle 111]$ and $[111|11\bar{2}\rangle\bar{1}\bar{1}\bar{1}]=[\bar{1}\bar{1}\bar{1}|\bar{1}\bar{1}2\rangle 111]$ walls corresponding to the segments of the the zigzag wall in Fig. 3b. All these three equivalent walls are realized in the three sides of the triangular prism in Fig. 4a.
The irreversible decomposition of DW2 wall documented in Fig. 3 and the systematic preference for DW1-type domain walls in stable prismatic nanodomains (Fig. 4a) suggests that the planar domain wall energy density $\sigma_{2}$ of DW2 is significantly higher then $\sigma_{1}$ of DW1.
The broadly applied straightforward procedure determination of the domain wall energy density from computational experiments consists in calculation of energy of planar domains with desired crystallographic orientation in a periodic cell like that of Fig. 1, subtracting the corresponding energy of the single domain state, and dividing the difference by the total surface of the walls within the supercell. In case of non-reciprocal wall pair, however, this yields only the average energy density $(\sigma_{1}+\sigma_{2})/2$. In order to overcome this topological obstacle, we have thus used the prismatic nanodomain of Fig. 4a. As argued above, the all three boundaries of these nanodomains are equivalent to DW1. For elimination of the domain fold energy contributions, we have tuned the initial conditions in a way to obtain a whole sequence of similar relaxed nanodomain configurations but with the base edge $l$ ranging from 5 to 20 nm. The slope of the excess energy associated with the domain walls vs the base edge lenght $l$ shown in Fig. 4b yielded $\sigma_{1}\approx 40$ mJ/m2. In combination with the average energy density obtained from the standard configuration of Fig. 1, we could see that $\sigma_{2}\approx 130$ mJ/m2, which is indeed significantly larger than $\sigma_{1}$.
In summary, we have explored the R180 domain walls within BiFeO3 and find out that so far little investigated ($\bar{1}\bar{1}2$)-oriented R180 walls are the lowest energy ones. We have given symmetry arguments why these walls are non-reciprocal and explained various profound consequences of this peculiar property.
This work was supported by the Czech Science Foundation (project no. 19-28594X) and by the European Union’s Horizon 2020 research and innovation programme under grant agreement no. 964931 (TSAR). MAPG was supported by the European Union and the Czech Ministry of Education, Youth and Sports (Project: MSCA Fellowship CZ FZU I - CZ.02.01.01/00/22__\__010/0002906).